# Barriers for Learning in an Evolving World: Mathematical Understanding of Loss of Plasticity

### v1

@@ Appendix A Theoretical Appendix
A.1 Formal proof of the rank–gain theorem under non-linear activations
This section states and proves in full detail the theorem used in Sec.˜3. We work with standard Gaussian inputs and use Hermite expansions to characterize the correlation kernel of a nonlinearity.
Hermite basis and notation
Let $(h_{k})_{k\geq 0}$ be the orthonormal probabilists’ Hermite polynomials in $L^{2}(\gamma)$ where $\gamma=\mathcal{N}(0,1)$, with $h_{0}(z)=1$, $h_{1}(z)=z$, and $\mathbb{E}[h_{k}(Z)h_{\ell}(Z)]=\delta_{k\ell}$ for $Z\sim\gamma$. Any $\phi\in L^{2}(\gamma)$ admits a Hermite expansion
@@ A.1 Formal proof of the rank–gain theorem under non-linear activations
Write $\sigma_{\phi}^{2}=\mathrm{Var}(\phi(Z))=\sum_{k\geq 1}a_{k}^{2}$. We call $\phi$ nonlinear when at least one coefficient with $k\geq 2$ is nonzero.
Correlation kernel
For jointly Gaussian $(X,Y)$ with $\mathbb{E}X=\mathbb{E}Y=0$, $\mathrm{Var}(X)=\mathrm{Var}(Y)=1$, and $\mathrm{corr}(X,Y)=r\in[-1,1]$, define
$K_{\phi}(r)\;=\;\mathrm{Corr}\big(\phi(X),\phi(Y)\big)\;=\;\frac{\mathbb{E}[\phi(X)\phi(Y)]-\mathbb{E}[\phi(X)]\mathbb{E}[\phi(Y)]}{\sqrt{\mathrm{Var}(\phi(X))\,\mathrm{Var}(\phi(Y))}}\,.$
Using the Hermite expansion and Mehler’s identity $\mathbb{E}[h_{k}(X)h_{\ell}(Y)]=\delta_{k\ell}\,r^{k}$ gives
$K_{\phi}(r)=\frac{\sum_{k\geq 1}a_{k}^{2}r^{k}}{\sum_{k\geq 1}a_{k}^{2}}=\sum_{k\geq 1}w_{k}r^{k},\qquad w_{k}:=\frac{a_{k}^{2}}{\sum_{\ell\geq 1}a_{\ell}^{2}}\ \geq 0,\ \ \sum_{k\geq 1}w_{k}=1.$ (1)
Hence $K_{\phi}$ is a convex combination of the monomials $r^{k}$ for $k\geq 1$.
Lemma A.1 (basic properties of $K_{\phi}$).
If $\phi$ is nonlinear with $\phi\in L^{2}(\gamma)$, then the correlation kernel $K_{\phi}$ defined in (1) satisfies
$K_{\phi}(0)=0,\qquad K_{\phi}(1)=1,\qquad|K_{\phi}(r)|<|r|\text{ for all }|r|<1.$
@@
The first two identities follow by evaluating (1) at $r=0$ and $r=1$. For the strict inequality, write $K_{\phi}(r)=\sum_{k\geq 1}w_{k}r^{k}$ with weights $w_{k}\geq 0$, $\sum_{k}w_{k}=1$. If $|r|<1$ and there exists $k\geq 2$ with $w_{k}>0$ (nonlinearity), then
$|K_{\phi}(r)|\leq\sum_{k\geq 1}w_{k}|r|^{k}=|r|\Big(w_{1}+\sum_{k\geq 2}w_{k}|r|^{k-1}\Big)<|r|\Big(w_{1}+\sum_{k\geq 2}w_{k}\Big)=|r|.$
Lemma A.2 (entrywise action on Gaussian correlation matrices).
Let $Z=(Z_{1},\dots,Z_{d})^{\top}\sim\mathcal{N}(0,C)$ with $C$ a correlation matrix. For $\phi\in L^{2}(\gamma)$,
$\mathrm{Corr}\big(\phi(Z_{i}),\phi(Z_{j})\big)\;=\;K_{\phi}(C_{ij})\qquad\text{for all }i,j.$
Equivalently, the post-activation correlation matrix equals $K_{\phi}(C)$ entrywise. Moreover,
$K_{\phi}(C)=\sum_{k\geq 1}w_{k}\,C^{\odot k},$
where $C^{\odot k}$ is the $k$-th Hadamard power and $(w_{k})_{k\geq 1}$ are as in (1). Hence $K_{\phi}(C)$ is a correlation matrix: it is positive semidefinite and has unit diagonal.
Proof.
Fix $i,j$. By the same Hermite calculation used for (1), with $r=C_{ij}$ we have
$\mathrm{Corr}\big(\phi(Z_{i}),\phi(Z_{j})\big)=\frac{\sum_{k\geq 1}a_{k}^{2}r^{k}}{\sum_{k\geq 1}a_{k}^{2}}=K_{\phi}(r).$
Stacking these equalities over all pairs $(i,j)$ yields the entrywise identity and the series $K_{\phi}(C)=\sum_{k\geq 1}w_{k}C^{\odot k}$. Each Hadamard power $C^{\odot k}$ is positive semidefinite by the Schur product theorem, and the diagonal entries are $(C_{ii})^{k}=1$, so the nonnegative convex combination is positive semidefinite with unit diagonal. ∎
Lemma A.3 (Frobenius contraction).
For any correlation matrix $C$ and nonlinear $\phi$ with kernel $K_{\phi}$ as above,
$\sum_{i\neq j}K_{\phi}(C_{ij})^{2}\;\leq\;\sum_{i\neq j}C_{ij}^{2},$
with strict inequality if there exists $i\neq j$ such that $|C_{ij}|<1$.
Proof.
By Lemma˜A.1, $|K_{\phi}(r)|<|r|$ for every $|r|<1$, and trivially $|K_{\phi}(r)|\leq|r|$ for $|r|\leq 1$. Apply this pointwise to each off-diagonal entry $C_{ij}$ and sum the squares. If some $|C_{ij}|<1$, the corresponding term is strictly reduced, and no term increases, so the sum is strictly reduced. ∎
We can now state and prove the theorem used in the main text.
Theorem A.1 (rank gain across one linear–nonlinear step).
Let $C$ be a $d\times d$ correlation matrix and let $\phi\in L^{2}(\gamma)$ be nonlinear. Define $K_{\phi}$ by (1). Then $K_{\phi}(C)$ is a correlation matrix and
$\frac{\mathrm{er}_{2}\big(K_{\phi}(C)\big)}{\mathrm{er}_{2}(C)}\;=\;\frac{d+\sum_{i\neq j}C_{ij}^{2}}{d+\sum_{i\neq j}K_{\phi}(C_{ij})^{2}}\;\geq\;1.$
Moreover, the ratio is strictly greater than $1$ whenever there exists $i\neq j$ with $|C_{ij}|<1$. If the ratio equals $1$, then every off-diagonal magnitude satisfies $|C_{ij}|\in\{0,1\}$.
Proof.
By Lemma˜A.2, $K_{\phi}(C)$ is a correlation matrix and $\mathrm{tr}\,K_{\phi}(C)=d=\mathrm{tr}\,C$. For any correlation matrix $M$,
$\|M\|_{F}^{2}\;=\;\sum_{i,j}M_{ij}^{2}\;=\;d+\sum_{i\neq j}M_{ij}^{2},$
hence
$\mathrm{er}_{2}(M)=\frac{(\mathrm{tr}\,M)^{2}}{\|M\|_{F}^{2}}=\frac{d^{2}}{d+\sum_{i\neq j}M_{ij}^{2}}.$
Applying this to $M=C$ and $M=K_{\phi}(C)$ yields the displayed ratio. Lemma˜A.3 gives $\sum_{i\neq j}K_{\phi}(C_{ij})^{2}\leq\sum_{i\neq j}C_{ij}^{2}$, which makes the ratio at least $1$, and strictly larger than $1$ if some $|C_{ij}|<1$.
For the equality case, suppose the ratio equals $1$. Then the two Frobenius norms coincide, so
$\sum_{i\neq j}\Big(C_{ij}^{2}-K_{\phi}(C_{ij})^{2}\Big)=0.$
Each term in the sum is nonnegative by Lemma˜A.1. Therefore every term must vanish, that is, for all $i\neq j$,
$K_{\phi}(C_{ij})^{2}=C_{ij}^{2}.$
If $|C_{ij}|<1$ and $\phi$ is nonlinear, Lemma˜A.1 gives $|K_{\phi}(C_{ij})|<|C_{ij}|$, a contradiction. Hence for all $i\neq j$ we must have $|C_{ij}|\in\{0,1\}$, which proves the final claim. ∎
Remark A.1 (dependence on operating regime).
If the pre-activation is reparameterized as $\phi_{a,b}(z)=\phi(az+b)$, the Hermite coefficients and thus the weights $(w_{k})$ in (1) change. The contraction at intermediate correlations can be summarized by the decorrelation strength
$\kappa_{\phi}\;=\;\max_{r\in[0,1]}\big(r^{2}-K_{\phi}(r)^{2}\big),$
which quantifies the maximal per-entry reduction of squared correlation. For instance, more negative effective bias for ReLU and larger effective gain for $\tanh$ increase $\kappa_{\phi}$, while also reducing typical derivatives, connecting rank gains to the emergence of frozen units.
A.1.1 Activation modulation increases decorrelation and also induces frozen units
Training changes pre–activation statistics, which modifies the operating regime of $\phi$. Parameterize $\phi_{a,b}(z)=\phi(az+b)$. Both the kernel $K_{\phi}$ and the decorrelation strength $\kappa_{\phi}$ vary with $(a,b)$. Section˜A.1 shows that regimes that raise decorrelation also reduce typical derivatives, yielding frozen units:
  • For $\tanh$, increasing gain $a$ raises the local decorrelation rate $\alpha_{\phi}=\mathbb{E}[\phi^{\prime}(Z)^{2}]/\mathbb{E}[\phi(Z)^{2}]-1$ while $\phi^{\prime}(az)\to 0$ for almost all $z$ as $a\to\infty$.
  • For ReLU, making the effective bias $b$ negative with $|b|/a$ large increases $\alpha_{\phi}$ while $\mathbb{P}(az+b<0)\to 1$, so $\phi^{\prime}(az+b)=0$ on most inputs.
A.2 LoP manifolds: Formal Statement and Proof
@@
This section provides the formal definitions, statement, and proof for the LoP manifold theorem. Since the frozen manifold argument is self explanatory, we will only prove the cloning manifold result that is more non-trivial. First, let us introduce our neural network network formalization. A feed-forward neural network is defined by a directed acyclic graph $G=(V,E,w)$, where $V$ is the set of nodes (neurons), $E$ is the set of directed edges (connections), $E$ representations the structure of the computational graph of the network, and $w:\mathbb{E}\to\mathbb{R},$ is the weight parameters of the network, which will be denoted $w(u,v)$ for each edge $(u,v)\in E$. Furthermore, $V_{\text{in}}\subset V$ are the input nodes, and $V_{\text{out}}\subset V$ are the output nodes. The post-activation $h(v)$ of a node $v\in V$ is computed as:
@@
    $\mathcal{M}_{RE}=\left\{w\in\mathbb{R}^{|E|}\;\middle|\;\forall(i,j)\in\widetilde{E},\text{ and }\forall r,r^{\prime}\in S_{i},\text{ it holds }\sum_{u\in S_{j}}w_{ru}=\sum_{u\in S_{j}}w_{r^{\prime}r}\right\}$
@@
Let $G=(V,E,W),$ denote a network that is partitioned with $S_{1},\dots,S_{k}.$ For any input and label $(x,y)$:
@@
    If $W\in\mathcal{M}_{RE}$, then all units in the same cluster $u,v\in S_{k}$ have identical forward activations $h(u)=h(v).$
@@
Proof of Theorem˜A.2 (Cloned-Unit Manifold).
@@
Row-equitability assumption implies identical forward for units in the same block. The induction hypothesis is that for all $k,$ any preceding unit $p\leq k,$ that belongs same partition $u_{p},u_{k}\in S_{i},$ will have identical forward $h(u_{p})=h(u_{k}).$ Because cloning does not apply to input units, meaning that every unit is a separate block, the hypothesis trivially holds for all input units $k=1,\dots,d$ where $d$ is input dimension. Now, let us prove the induction step, assuming step $k.$ Let $p\leq k$ correspond to a unit in the same block $u_{p},u_{k}\in S_{i}.$ Now, consider all the units that have incoming edges to these two units, which necessarily must appear before $p$. Let’s consider all such units within the same block $S_{j}.$ Because these units appear before $k,$ the induction hypothesis tells us that they have identical forward. Thus, the total contribution from these units to pre-activations $z(u_{p})$ and $z(u_{k})$ will be proportional to sum of edge weights from units in $S_{j}$. Because of our construction of the ordering, all the units in $S_{j}$ that feed into $S_{i}$ must occur before them. Now, the row-equal assumption implies that the sum of weights from all these units to $u_{p}$ and $u_{k}$ must be equal weight sum. Thus, we have proven that pre-activation contribution from units in $S_{j}$ will be identical for $u_{p}$ and $u_{k}$. Because we chose $S_{j}$ arbitrarily and it could have been any block, we have proven that pre-activation these units must be identical $z(u_{p})=z(u_{k})$. Since they also have identical activation function, they will have identical outputs $h(u_{p})=h(u_{k})$. This completes the induction hypothesis for forward pass cloning.
@@
Backward cloning. We want to prove that column and row-equitability assumption implies identical backward for units in the same block. The proof strategy will be highly similar to the forward cloning case, with the key difference that our induction will be backward in our ordering, starting from latest output units and then moving in backward in the list. The induction hypothesis for step $k$ is that, for al $q>k$, if they are in the same block $u_{k},u_{q}\in S_{i}$, they will have identical backwards $\delta(u_{k})=\delta(u_{q}).$ Because output units are not themselves cloned, the induction step holds trivially for the last output nodes. Now let us prove the induction hypothesis for $k$ assuming that it holds for all higher steps. Now, for some arbitrary block $S_{j}$ that units in $S_{i}$ feed into, consider all outgoing connections from $u_{k},u_{q}$ to the units in this block. Because of our construction of the ordering, all the units in $S_{j}$ that $S_{i}$ feeds into must occur after $S_{j}.$ Thus, by induction hypothesis, all these units must have identical backwards. Furthermore, from our column-equitability assumption we know that total edge weights from $u_{k},u_{q}$ to these units must be identical. Thus, the summation formulas in the backward of $u_{k}$ and $u_{q}$ are similar. Finally, since $S_{j}$ was chosen arbitrarily, this summation is identical for all subsequent blocks, which implies the overal sum is also identical. To conclude the proof, note that because of row-equitability condition we already inherit the proof from the forward case, implying $f^{\prime}(z(u_{k}))=f^{\prime}(z(u_{q})).$ Thus, both parts to the backward formula for $u_{k},u_{q}$ will be identical, which proves they have identical backwards. This finishes the induction step.
@@
A.3 Modular cloning profiles and a composition theorem
This subsection formalizes a modular extension of the cloning theorem in Theorem˜2.1 (see also Sec.˜A.2) and proves that local, module-level cloning certificates glue to yield cloning for the entire composed architecture.

### v2

@@ Appendix A Theoretical Appendix
A.1 Formal proof of the Rank Gain and Decorrelation Theorem
This section provides the rigorous derivation of Theorem 3.1, establishing a quantitative lower bound on the increase in Rényi-2 effective rank across a nonlinear layer.
Preliminaries: Hermite Expansion and Kernel Structure
We consider an activation function $\phi\in L^{2}(\gamma)$, where $\gamma=\mathcal{N}(0,1)$. Let $\{h_{k}\}_{k\geq 0}$ be the orthonormal Hermite basis. The activation admits the expansion:
@@ A.1 Formal proof of the Rank Gain and Decorrelation Theorem
We follow the assumptions of Theorem 3.1, where $\phi$ is standardized such that $\mathbb{E}[\phi(Z)]=0$ (implying $a_{0}=0$) and $\mathrm{Var}(\phi(Z))=1$ (implying $\sum_{k\geq 1}a_{k}^{2}=1$).
The correlation kernel $K_{\phi}(r)$ is defined for jointly Gaussian $(X,Y)$ with correlation $r$:
$K_{\phi}(r)=\mathrm{Corr}\big(\phi(X),\phi(Y)\big).$
Using Mehler’s identity, the kernel is given by:
$K_{\phi}(r)=\sum_{k=1}^{\infty}a_{k}^{2}r^{k}=\sum_{k=1}^{\infty}w_{k}r^{k},$ (1)
where $w_{k}=a_{k}^{2}$. Key properties derived from this structure are:
  1. 1.
    $w_{k}\geq 0$ for all $k$.
  2. 2.
    $\sum_{k=1}^{\infty}w_{k}=1$, thus $K_{\phi}(1)=1$ and $K_{\phi}(0)=0$.
  3. 3.
    The linearity coefficient is $\alpha_{\phi}=K_{\phi}^{\prime}(0)=w_{1}$. If $\phi$ is nonlinear, $\alpha_{\phi}<1$.
  4. 4.
    $|K_{\phi}(r)|\leq|r|$ for $r\in[-1,1]$.
The Universal Quartic Gap Bound
The key to the quantitative rank gain theorem is establishing a rigorous lower bound on the decorrelation gap $G(r)=r^{2}-K_{\phi}(r)^{2}$.
Lemma A.1 (Universal Quartic Gap Bound).
Let $\alpha=K_{\phi}^{\prime}(0)$. The decorrelation strength is defined as $\gamma_{\phi}=\frac{1-\alpha}{1+\alpha}$. For any correlation $r\in[-1,1]$:
$r^{2}-K_{\phi}(r)^{2}\;\geq\;\gamma_{\phi}\cdot r^{2}(1-r^{2}).$ (2)
@@
We aim to find the kernel that maximizes $|K_{\phi}(r)|$ for a fixed $\alpha$, subject to the constraints $w_{k}\geq 0$ and $\sum w_{k}=1$.
Let $x=|r|\in[0,1]$. Due to the non-negativity of $w_{k}$:
$|K_{\phi}(r)|=|\sum_{k=1}^{\infty}w_{k}r^{k}|\leq\sum_{k=1}^{\infty}w_{k}|r|^{k}=K_{\phi}(x).$
We now seek an upper bound for $K_{\phi}(x)$. Since $x\in[0,1]$, we have $x^{k}\leq x^{2}$ for all $k\geq 2$.
$K_{\phi}(x)$ $=w_{1}x+\sum_{k=2}^{\infty}w_{k}x^{k}$
$\leq\alpha x+x^{2}\sum_{k=2}^{\infty}w_{k}$
$=\alpha x+x^{2}(1-\alpha).$
Let $Q(x)=\alpha x+(1-\alpha)x^{2}$. This quadratic function is the extremal kernel that maximizes the magnitude for a given $\alpha$ under these constraints.
Now we bound the gap $G(r)$:
$G(r)=r^{2}-K_{\phi}(r)^{2}\geq x^{2}-Q(x)^{2}.$
We must verify that $x^{2}-Q(x)^{2}\geq\gamma_{\phi}x^{2}(1-x^{2})$.
$x^{2}-[\alpha x+(1-\alpha)x^{2}]^{2}\geq\frac{1-\alpha}{1+\alpha}x^{2}(1-x^{2}).$
For $x\in(0,1)$, we divide both sides by the positive term $x^{2}(1-x)$:
$\frac{1-[\alpha+(1-\alpha)x]^{2}}{1-x}\geq\frac{1-\alpha}{1+\alpha}(1+x).$
We factor the LHS numerator as a difference of squares $1-Y^{2}=(1-Y)(1+Y)$, where $Y=\alpha+(1-\alpha)x$. Note that $1-Y=1-\alpha-(1-\alpha)x=(1-\alpha)(1-x)$. The LHS simplifies to:
$\text{LHS}=\frac{(1-\alpha)(1-x)[1+\alpha+(1-\alpha)x]}{1-x}=(1-\alpha)[1+\alpha+x(1-\alpha)].$
The inequality becomes:
$(1-\alpha)[1+\alpha+x(1-\alpha)]\geq\frac{1-\alpha}{1+\alpha}(1+x).$
Assuming $\phi$ is nonlinear ($\alpha<1$), we can divide by $(1-\alpha)$. We then multiply by $(1+\alpha)$:
$(1+\alpha)[1+\alpha+x(1-\alpha)]\geq 1+x.$
$(1+\alpha)^{2}+x(1-\alpha)(1+\alpha)\geq 1+x.$
$1+2\alpha+\alpha^{2}+x(1-\alpha^{2})\geq 1+x.$
$1+2\alpha+\alpha^{2}+x-\alpha^{2}x\geq 1+x.$
$2\alpha+\alpha^{2}(1-x)\geq 0.$
Since $\alpha\geq 0$ (as $w_{1}=a_{1}^{2}$) and $x\in[0,1]$, the inequality holds true. The boundary cases $r=0,r=\pm 1,\alpha=1$ hold trivially. ∎
Proof that rank increases under non-linearity
We now connect the gap bound (Lemma A.1) to the increase in Rényi-2 rank.
Proof of Theorem 3.1 (Rank Gain and Decorrelation Potential).
Let $C$ be a $d\times d$ correlation matrix. The Rényi-2 effective rank is $\mathrm{er}_{2}(C)=(\text{tr}\,C)^{2}/\|C\|_{F}^{2}$. Since $C$ and $K_{\phi}(C)$ are correlation matrices, $\text{tr}(C)=\text{tr}(K_{\phi}(C))=d$.
1. Rank Ratio: The ratio of the effective ranks is:
$R(C)=\frac{\mathrm{er}_{2}(K_{\phi}(C))}{\mathrm{er}_{2}(C)}=\frac{d^{2}/\|K_{\phi}(C)\|_{F}^{2}}{d^{2}/\|C\|_{F}^{2}}=\frac{\|C\|_{F}^{2}}{\|K_{\phi}(C)\|_{F}^{2}}.$
2. Total Decorrelation Gap $\Delta(C)$: We define the total reduction in the squared Frobenius norm:
$\Delta(C)=\|C\|_{F}^{2}-\|K_{\phi}(C)\|_{F}^{2}.$
Since $\|M\|_{F}^{2}=d+\sum_{i\neq j}M_{ij}^{2}$ for a correlation matrix $M$:
$\Delta(C)=\sum_{i\neq j}(C_{ij}^{2}-K_{\phi}(C_{ij})^{2}).$
3. Applying the Bound: We apply the Universal Quartic Gap Bound (Lemma A.1) to each term:
$\Delta(C)\geq\sum_{i\neq j}\gamma_{\phi}C_{ij}^{2}(1-C_{ij}^{2})=\gamma_{\phi}\Psi(C).$
4. Bounding the Rank Ratio: We express the rank ratio using the gap $\Delta(C)$:
$R(C)=\frac{\|C\|_{F}^{2}}{\|C\|_{F}^{2}-\Delta(C)}.$
Substituting the lower bound for $\Delta(C)$:
$R(C)\geq\frac{\|C\|_{F}^{2}}{\|C\|_{F}^{2}-\gamma_{\phi}\Psi(C)}=\frac{1}{1-\gamma_{\phi}\frac{\Psi(C)}{\|C\|_{F}^{2}}}.$
We use the inequality $(1-u)^{-1}\geq 1+u$, valid for $u\in[0,1)$. Let $u=\gamma_{\phi}\frac{\Psi(C)}{\|C\|_{F}^{2}}$. We know $u\geq 0$. We verify $u<1$. Since $\Delta(C)\geq\gamma_{\phi}\Psi(C)$, we have $u\leq\frac{\Delta(C)}{\|C\|_{F}^{2}}$. As $\|K_{\phi}(C)\|_{F}^{2}\geq d>0$, we have $\Delta(C)<\|C\|_{F}^{2}$. Thus $u<1$.
Applying the inequality:
$R(C)\geq 1+\gamma_{\phi}\frac{\Psi(C)}{\|C\|_{F}^{2}}.$
5. Equality Conditions: The rank ratio equals 1 if and only if $\Delta(C)=0$. This implies $\gamma_{\phi}\Psi(C)=0$. This occurs if $\gamma_{\phi}=0$ (the activation is linear, $\alpha=1$) or if $\Psi(C)=0$ (all correlations $C_{ij}\in\{0,\pm 1\}$). ∎
Validation of the Quartic Gap Bound.
Figure A.1 (Top Row) validates the Universal Quartic Gap Bound (Lemma˜A.1). For representative activations, we plot the actual decorrelation gap $\Delta(c)=c^{2}-K_{\phi}(c)^{2}$ against the theoretical lower bound $\gamma_{\phi}c^{2}(1-c^{2})$ over the correlation domain $c\in[-1,1]$.In all cases, the actual gap strictly dominates the lower bound. The gap vanishes only at the fixed points $c\in\{0,\pm 1\}$ or for linear functions ($\gamma_{\phi}=0$), confirming that nonlinear activations strictly reduce correlation magnitude (and thus increase rank) for any non-trivial input correlation structure.
Validation of Rank Gain.
Figure A.1 (Bottom Row) tests the Rank Gain Theorem (Theorem˜3.1). We generate a continuum of correlation matrices $C(\lambda)$ ranging from the identity matrix to a dense random matrix. We compute the actual Rényi-2 rank expansion ratio $R(C)=\mathrm{er}_{2}(K_{\phi}(C))/\mathrm{er}_{2}(C)$ and compare it to the derived lower bound $1+\gamma_{\phi}\Psi(C)/\|C\|_{F}^{2}$.The results show that the actual rank gain consistently exceeds the theoretical guarantee. The gain is positive whenever the decorrelation potential $\Psi(C)$ is non-zero, demonstrating that the nonlinearity actively consumes correlation structure to generate effective rank, with an efficiency bounded by $\gamma_{\phi}$.
Figure A.1: Decorrelation and rank-gain under non-linearity. Empirical analysis activations theoretical predictions about quadratic lower bound for decorrelatioin, and for rank gain potential.
A.2 Rank-plasticity tension: an empirical validation
This section shows that maximizing the de-correlation strength $\gamma_{\phi}$, necessitates driving the activation function into regimes where the gradient vanishes (freezing).
In particular, we consider modulating activations by scale $a$ and bias $b$: $\psi(z)=\phi(az+b)$, and analyze the de-correlation strength $\gamma(a,b)=\frac{(\mathbb{E}[\psi^{\prime}(Z)])^{2}}{\text{Var}(\psi(Z))}$ as a function of shape parameters $\{a,b\}$.
Note that $\gamma(a,b)=\bigl(\mathbb{E}[\psi^{\prime}(Z)]\bigr)^{2}/\operatorname{Var}(\psi(Z))$ coincides with the derivative of the corresponding correlation kernel at 0, i.e. $\gamma(a,b)=K^{\prime}_{\psi}(0)$ after normalization. Thus, $\gamma(a,b)$ and the $\gamma_{\phi}$ used in Theorem 3.1 are related by a simple monotone reparameterization.
Here we empirically validate that the bias $b$ for ReLU and scale $a$ for Tanh under standard Gaussian input to measure the trade-off between decorrelation and gradient availability. As shown in Figure A.2, the decorrelation coefficient $\gamma_{\phi}$ is strictly monotonic with respect to the modulation parameters for both activations. Crucially, the analysis reveals a severe rank-plasticity tension: approaching maximum decorrelation ($\gamma_{\phi}\to 1$) necessitates driving the activation into a “frozen” regime (dead zone for ReLU or saturation for Tanh) with probability approaching $1.0$, thereby effectively eliminating the gradient flow required for training.
To extend this beyond ReLU and Tanh, Figure A.3 visualizes the fundamental trade-off described in Sec.˜3. For all activations we plot the decorrelation strength $\gamma_{\phi}$ against the probability that the unit is frozen ($P(|f^{\prime}|<\epsilon)$ where $\epsilon=0.01$). The results confirm that to achieve high rank expansion capabilities (large $\gamma_{\phi}$), in most cases the parameters must push the activation into a regime where the gradient vanishes for the vast majority of inputs, demonstrating the tension between feature learning and trainability.
Figure A.2: The rank-plasticity tradeoff for ReLU and Tanh. Empirical analysis of ReLU and Tanh activations confirms that increasing the decorrelation coefficient $\gamma_{\phi}$ (to improve rank) strictly requires increasing the probability that the unit is frozen (dead or saturated), approaching $1.0$ as $\gamma_{\phi}\to 1$.
Figure A.3: The rank-plasticity tradeoff for general activations. Empirical analysis of all activations $\gamma_{\phi}$ vs the probability that the unit is frozen, when we vary the pre-activations mean and bias.
A.3 LoP manifolds: Formal Statement and Proof
@@
This section provides the formal definitions, statement, and proof for the LoP manifold theorem. Since the frozen manifold argument is self explanatory, we will only prove the cloning manifold result that is more non-trivial. First, let us introduce our neural network formalization. A feed-forward neural network is defined by a directed acyclic graph $G=(V,E,w)$, where $V$ is the set of nodes (neurons), $E$ is the set of directed edges (connections), $E$ representations the structure of the computational graph of the network, and $w:\mathbb{E}\to\mathbb{R},$ is the weight parameters of the network, which will be denoted $w(u,v)$ for each edge $(u,v)\in E$. Furthermore, $V_{\text{in}}\subset V$ are the input nodes, and $V_{\text{out}}\subset V$ are the output nodes. The post-activation $h(v)$ of a node $v\in V$ is computed as:
@@
    $\mathcal{M}_{RE}=\left\{w\in\mathbb{R}^{|E|}\;\middle|\;\forall(i,j)\in\widetilde{E},\text{ and }\forall r,r^{\prime}\in S_{i},\text{ it holds }\sum_{u\in S_{j}}w_{ru}=\sum_{u\in S_{j}}w_{r^{\prime}u}\right\}$
@@
Let $G=(V,E,W),$ denote a network that is partitioned with $S_{1},\dots,S_{k}.$ Furthermore, units within a single block share the same activation function $\forall u,v\in S_{k}:f_{v}=f_{u},$ and there are no edges within a block. For any input and label $(x,y)$:
@@
    If $W\in\mathcal{M}_{CE}$, then all units in the same cluster $u,v\in S_{k}$ have identical forward activations $h(u)=h(v).$
@@
Proof of Theorem˜A.1 (Cloned-Unit Manifold).
@@
Column-equitability assumption implies identical forward for units in the same block. The induction hypothesis is that for all $k,$ any preceding unit $p\leq k,$ that belongs same partition $u_{p},u_{k}\in S_{i},$ will have identical forward $h(u_{p})=h(u_{k}).$ Because cloning does not apply to input units, meaning that every unit is a separate block, the hypothesis trivially holds for all input units $k=1,\dots,d$ where $d$ is input dimension. Now, let us prove the induction step, assuming step $k.$ Let $p\leq k$ correspond to a unit in the same block $u_{p},u_{k}\in S_{i}.$ Now, consider all the units that have incoming edges to these two units, which necessarily must appear before $p$. Let’s consider all such units within the same block $S_{j}.$ Because these units appear before $k,$ the induction hypothesis tells us that they have identical forward. Thus, the total contribution from these units to pre-activations $z(u_{p})$ and $z(u_{k})$ will be proportional to sum of edge weights from units in $S_{j}$. Because of our construction of the ordering, all the units in $S_{j}$ that feed into $S_{i}$ must occur before them. Now, the column-equal assumption implies that the sum of weights from all these units to $u_{p}$ and $u_{k}$ must be equal weight sum. Thus, we have proven that pre-activation contribution from units in $S_{j}$ will be identical for $u_{p}$ and $u_{k}$. Because we chose $S_{j}$ arbitrarily and it could have been any block, we have proven that pre-activation these units must be identical $z(u_{p})=z(u_{k})$. Since they also have identical activation function, they will have identical outputs $h(u_{p})=h(u_{k})$. This completes the induction hypothesis for forward pass cloning.
@@
Backward cloning. We want to prove that column and row-equitability assumption implies identical backward for units in the same block. The proof strategy will be highly similar to the forward cloning case, with the key difference that our induction will be backward in our ordering, starting from latest output units and then moving in backward in the list. The induction hypothesis for step $k$ is that, for al $q>k$, if they are in the same block $u_{k},u_{q}\in S_{i}$, they will have identical backwards $\delta(u_{k})=\delta(u_{q}).$ Because output units are not themselves cloned, the induction step holds trivially for the last output nodes. Now let us prove the induction hypothesis for $k$ assuming that it holds for all higher steps. Now, for some arbitrary block $S_{j}$ that units in $S_{i}$ feed into, consider all outgoing connections from $u_{k},u_{q}$ to the units in this block. Because of our construction of the ordering, all the units in $S_{j}$ that $S_{i}$ feeds into must occur after $S_{j}.$ Thus, by induction hypothesis, all these units must have identical backwards. Furthermore, from our row-equitability assumption we know that total edge weights from $u_{k},u_{q}$ to these units must be identical. Thus, the summation formulas in the backward of $u_{k}$ and $u_{q}$ are similar. Finally, since $S_{j}$ was chosen arbitrarily, this summation is identical for all subsequent blocks, which implies the overal sum is also identical. To conclude the proof, note that because of row-equitability condition we already inherit the proof from the forward case, implying $f^{\prime}(z(u_{k}))=f^{\prime}(z(u_{q})).$ Thus, both parts to the backward formula for $u_{k},u_{q}$ will be identical, which proves they have identical backwards. This finishes the induction step.
@@
A.4 Modular cloning profiles and a composition theorem
This subsection formalizes a modular extension of the cloning theorem in Theorem˜2.1 (see also Sec.˜A.3) and proves that local, module-level cloning certificates glue to yield cloning for the entire composed architecture.